Juan has spent $ 2,700 on furniture and electronics to fix his first apartment. He must pay $ 500 at the beginning and agree to pay the balance at a rate of 7% for 2 years.
a. Determine the total amount you will be financing.
b. Find the total interest.
C. Determine the amount owed.
d. Determine the monthly payment.
Answer #1

a. Total amount you will be financing=

2700-500= 2200

b.The total interest

The present value of the loan= 2200
first, we need to find the monthly payment , to determine the total interest amt.
Using the Formula to find PV of a loan,
PV=Pmt.*(1-(1+r)^-n)/r
where, PV = the present value of the loan= 2200
Pmt.= the monthly pmt. To be found out--- ??
r= 7% p.a. ie. 7%/12= 0.583% or 0.00583 p.m
n= no.of mthly. Pmts.= 2yrs.*12=24 mths
Now, plugging in the values,in the formula,
2200=Pmt.*(1-(1+0.00583)^-24)/0.00583
Solving the above, we get the mthly. Pmt. On the loan as
2200/((1-(1+0.00583)^-24)/0.00583)=
98.50
So, the total interest on the loan =
Total of the Mthly.pmts-the principal loan amt.
ie.(98.50*24)-2200=
164
c. Amt. owed=
PV of loan + initial pmt.=2200+500=2700
d. Mthly pmt.= $ 98.50
(as in b. above)